OpenStudy (anonymous):

coefficient of x^6 ( 2x^3- 1/x^2) ^7

OpenStudy (campbell_st):

the term in x^6 occurs when r = 3 \[^{7}C_{3} (2x^3)^{7 - 3} \times (-\frac{1}{x^2})^3\] just evaluate

OpenStudy (anonymous):

its says find the coefficient in the expansion . so its that the answer ?

OpenStudy (campbell_st):

well all you need to do is the expansion....

OpenStudy (anonymous):

okay good :) thx !

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

it do , just joined this group today